# Conformally invariant boundary arcs in double dimers

### v1

@@ 5. Loops and arcs on shifted domains
Loops get a weight $2$, arcs enclosing $z$ get a weight $21\pm\alpha/$ 2 according to their orientation, so we obtain, using the same notations as in the preceding proof,
@@
We use the same notations as in the preceding proof: $E_{\varepsilon}$ for the zipper, $\text{sgn}(E_{\varepsilon})_{w,b}$ for the direction of an edge of the zipper. For $i\in\{1,2\}$, we write the vertices of $\mathcal{G}_{\varepsilon}^{i}$ as $(x,i)$ with $x\in\mathcal{G}_{\varepsilon}$ if $i=1$, $x\in\mathcal{G}_{\varepsilon}\setminus\mathbb{R}\times\{0\}$ if $i=2$. Contrary to the last section, the matrix $K$ is a block matrix since $\psi=\begin{pmatrix}1\\ 0\end{pmatrix}$: $K_{(w,i),(b,j)}=0$ as soon as $i\neq j$. On the $i=j=1$ block, it is the Kasteleyn matrix on the Temperleyan domain $\mathcal{G}_{\varepsilon}^{1}$: it has Dirichlet boundary conditions everywhere. We write $K_{(w,1),(b,1)}=K^{1}_{w,b}$. On the $i=j=2$ block, it is the Kasteleyn matrix on the piecewise Temperleyan domain $\mathcal{G}_{\varepsilon}^{2}$: it has Dirichlet boundary conditions everywhere, except on the horizontal axis where it has Neumann boundary conditions (see [20] for a general discussion of mixed boundary conditions for Kasteleyn matrices). We write $K_{(w,2),(b,2)}=K^{2}_{w,b}$. The asymptotic of $(K^{1})^{-1}$ and $(K^{2})^{-1}$ when $\varepsilon\to 0$ were obtained in [20] and are recalled in Lemma C.3. Note that contrary to the last section, the vertical orientation is correct for $K^{1}$ and $K^{2}$, hence we do not need any gauge change. We can write
@@
$S_{(w,i),(b,j)}=\mathbbm{1}\{w\sim b\}\zeta_{w,b}\text{sgn}(E_{\varepsilon})_{w,b}(-1)^{i-1}.$ (35)
@@
×$\times\left((K^{j_{i+1}})^{-1}_{b_{i},w_{i+1}}-\Delta(w_{i+1},b_{i})\right)\bigg{\}}.$
@@
after reindexing $r_{i+1}=r_{i}$. Now similarly to the above, if for some $i=0,\dots,n-1$, we have that $1-\kappa_{i+1}+\kappa_{i}+\tfrac{1}{2}(1-\tau_{i})$ is even (contrary to the last section), then the sum over $r_{i}=\pm 1$ gives zero. Hence the only terms that remain have for all $i=0,\dots,n-1$ that $1-\kappa_{i+1}+\kappa_{i}+\tfrac{1}{2}(1-\tau_{i})$ is odd, which is the same as $\mathbbm{1}\{\kappa_{i}\neq\kappa_{i+1}\}\neq\mathbbm{1}\{\tau=1\}$, which is the same as $\tau_{i}=(-1)^{\kappa_{i}}(-1)^{\kappa_{i+1}}$. The sum is then independent of $r$, and we obtain
@@
6. The case of a rectangle: computing the number of arcs crossing in the vertical direction
In this section, we consider a slightly different setting (strongly inspired by section 9 of [17]) and show that in some cases we can compute new interesting quantities by diagonalizing the Kasteleyn matrix. Let $G$ be the subgraph of $\mathbb{Z}^{2}$ with vertex set $[0,n]\times[1,m]$. We define its boundary to be =$\partial=\{0,n\}\times[1,m]$. We define a graph $\mathcal{G}^{r}$ by taking two copies of $G$ with vertices labeled as $(x,y,1)$ and $(x,y,2)$ for $(x,y)\in G$ and by identifying $(x,y,1)=(x,y,2)$ for $(x,y)\in\partial G$. In other words, $\mathcal{G}^{r}$ is the cylinder $\mathbb{Z}/(2n\mathbb{Z})\times[1,m]$ if we identify $(x,y,1)=(x,y)$ and $(x,y,2)=(-x,y)$, with the slight abuse of notation that we identify $x\in\mathbb{Z}$ with its projection $\bar{x}\in\mathbb{Z}/(2n\mathbb{Z})$ so that in particular, for all $y$, $(n,y,1)=(n,y)=(-n,y)=(n,y,2)$. A uniformly random dimer configuration on $\mathcal{G}^{r}$ gives a random configuration of disjoint loops, doubled edges and arcs, obtained by superimposing the two layers. Arcs link two boundary vertices that can be on the same side or on opposite sides of the cylinder.
@@ 6. The case of a rectangle: computing the number of arcs crossing in the vertical direction
Note that it is an example where $\partial$ has two connected components. It complements the strip Example 3.7 where it has one connected component.
@@
We are left with computing the left-hand side, or rather its square $\frac{\det(K_{a})}{\det{K}}$. We define the matrix $\bar{K}_{a}$ by $\bar{K}_{a}(w,b)=K_{a}(w,b)$ for all $b$ black and $w$ white in $\mathcal{G}^{r}$: it is the symmetric version of $K_{a}$ (which is by definition antisymmetric). Using the block structure of $K_{a}$ and $\bar{K}_{a}$, we get $\det(\bar{K}_{a})=-\det(K_{a})$ so the ratio of determinants is left unchanged. We now observe that the matrix $\bar{K}_{a}$ that we have constructed is exactly that of Kenyon. Writing $\lambda=\rho^{2}$ (which satisfies $a^{2n}=\lambda$, $X=\rho^{2}+\frac{1}{\rho^{2}}$), and since $K$ corresponds to $a=1$ and $X=2$, if $m$ is even, the results of Section 9 of Kenyon (conformal invariance of domino tiling) imply
@@
For $\rho\in\mathbb{C}$, such that $Y=\frac{\rho+\frac{1}{\rho}}{2}>0$, $Y^{2}=\frac{X+2}{4}$ so $X>-2$ and all (squared) factors on both sides are non-negative so we can take the square root:
@@ Appendix A Computation in the continuum (by Avelio Sepúlveda)
The objective of this section is to compute the expected values of the random variables $n(z)$ and $o(z)$ for the ALE. These random variables are clearly conformally invariant, as the ALE itself is. More precisely we plan to show the following.
@@
Let $\mathcal{A}$ be an ALE in the strip ×$\mathbb{R}\times[0,\pi/2]$, $n(z)$ the amount of arcs of $\mathcal{A}$ that separate the upper boundary from $z$ and $O(z)=(-1)^{n}(z)$. We have that
@@
Note importantly that these moments exactly coincide with Example 3.7.
@@
In what follows $\lambda=\sqrt{\pi/8}$. Let $D\subseteq\mathbb{H}$ be a domain with $\partial D=\partial_{1}\cup\partial_{2}$, $\partial_{1}\cap\partial_{2}=\emptyset$ and $\partial_{2}$ an interval of $\mathbb{R}$. We also define $\check{D}$ to be the domain obtained by reflecting $D$ on $\mathbb{H}$. We take $\Phi$ to be a GFF in a domain $D$ with boundary condition $\lambda$ on $\partial_{1}$ and 00 on $\partial_{2}$, and $\check{\Phi}$ be a GFF in $D$ with boundary condition 00 on $\partial_{1}$ and free on $\partial_{2}$. We couple $\Phi$ and $\check{\Phi}$ such that $\mathbb{A}_{-\lambda,\lambda}$ of $\Phi$ coincides with $\mathcal{A}$ – the ALE of $\check{\Phi}$ as in [19] (See Figure 14). Furthermore, we denote by $u(z)$ the harmonic function in $D$ with boundary values $\lambda\mathbf{1}_{\partial_{1}}$ so that $\Phi-u(z)$ is a GFF with 00 boundary conditions in $D$.
Figure 14. Representation of $\mathbb{A}_{-\lambda,\lambda}$ and $A$ in $S^{+}$. The harmonic function associated to $\Phi_{\mathbb{A}_{-\lambda,\lambda}}$ takes alternating values in ±$\pm 2\lambda$, on the other hand $\check{\Phi}_{A}$ starts with value 00 on top and each time it crosses an curve changes values of ±$\pm 2\lambda$ independently in each connected component.
@@
Let $\Phi_{\epsilon}$ and $\check{\Phi}_{\epsilon}$ denote the $\epsilon$-circle average of $\Phi$ and $\check{\Phi}$ respectively. A straightforward computation yields that (see for example Theorem 1.23 of [9])
$\mathbb{E}\left[(\Phi_{\epsilon}(z)-u(z))^{2}\right]-\frac{1}{2\pi}\log(1/\epsilon)\to\frac{1}{2\pi}\log(CR(z;D))\ \ \ \text{ as }\epsilon\to 0.$ (42)
Furthermore, as $\Gamma=\frac{1}{\sqrt{2}}(\Phi-u+\check{\Phi})$ has the law of a GFF with 00 boundary conditions on $\check{D}$ restricted to $D$, we have that
$\mathbb{E}\left[(\check{\Phi}_{\epsilon}(z))^{2}\right]-\frac{1}{2\pi}\log(1/\epsilon)\to\frac{1}{\pi}\log(CR(z;\check{D}))-\frac{1}{2\pi}\log(CR(z;D))\ \ \ \text{ as }\epsilon\to 0.$
@@
$\Phi=\Phi_{\mathbb{A}_{-\lambda,\lambda}}+\Phi^{\mathbb{A}_{-\lambda,\lambda}}\text{ and }\check{\Phi}=\check{\Phi}_{\mathcal{A}}+\check{\Phi}^{\mathcal{A}},$
where $\Phi_{\mathbb{A}_{-\lambda,\lambda}}$ and $\check{\Phi}_{\mathcal{A}}$ can be represented by harmonic functions thanks to [22], and conditionally on $\mathcal{A}$, $\Phi^{\mathbb{A}_{-\lambda,\lambda}}$ and $\check{\Phi}^{\mathcal{A}}$ are 0 boundary GFF on $D\backslash\mathcal{A}$ (conditionally) independent of $(\Phi_{\mathbb{A}_{-\lambda,\lambda}},\check{\Phi}_{\mathcal{A}})$. Furthermore, it is possible to check that
$\Phi_{\mathbb{A}_{-\lambda,\lambda}}(z)=\lambda O(z)\text{ and }\check{\Phi}_{\mathcal{A}}=\sum_{k=1}^{n(z)}\xi_{k},$
@@
$u^{2}(z)=\mathbb{E}\left[(\Phi_{\epsilon}(z))^{2}\right]=\mathbb{E}\left[((\Phi_{\mathbb{A}_{-\lambda,\lambda}})^{D}_{\epsilon}(z))^{2}\right]+\mathbb{E}\left[((\Phi^{\mathbb{A}_{-\lambda,\lambda}})_{\epsilon}(z))^{2}\right].$
@@
$u^{2}(z)$ $=\mathbb{E}\left[((\Phi_{\mathbb{A}_{-\lambda,\lambda}})_{\epsilon}(z))^{2}\right]+\mathbb{E}\left[((\Phi^{\mathbb{A}_{-\lambda,\lambda}})_{\epsilon}(z))^{2}-(\Phi_{\epsilon}(z)-u(z))^{2}\right]$
$\to\lambda^{2}+\frac{1}{2\pi}\mathbb{E}\left[\log\left(\frac{CR(z;D\backslash\mathbb{A}_{-2\lambda,2\lambda})}{CR(z;D)}\right)\right]\ \ \text{as }\epsilon\to 0.$ (43)
@@
$\mathbb{E}[(\check{\Phi}_{\epsilon}(z))^{2}]=\mathbb{E}\left[((\check{\Phi}_{\mathcal{A}})_{\epsilon}(z))^{2}\right]+\mathbb{E}\left[((\check{\Phi}^{\mathcal{A}})_{\epsilon}(z))^{2}\right].$ (44)
@@
00 0 $=\mathbb{E}\left[((\check{\Phi}_{\mathcal{A}})_{\epsilon}(z))^{2}\right]+\mathbb{E}\left[((\check{\Phi}^{\mathcal{A}})_{\epsilon}(z))^{2}-(\check{\Phi}_{\epsilon}(z))^{2}\right]$
$\to\mathbb{E}\left[\left(u(z)+2\lambda\sum_{k=1}^{n(z)}\xi_{k}\right)^{2}\right]+\frac{1}{2\pi}\mathbb{E}\left[-2\log(CR(z;\check{D}))+\log(CR(z;D))+\log\left(CR(z;D\backslash\mathcal{A})\right)\right]$
$=u(z)+4\lambda^{2}\mathbb{E}\left[n(z)\right]+\frac{1}{2\pi}\mathbb{E}\left[-2\log(CR(z;\check{D}))+\log(CR(z;D))+\log\left(CR(z;D\backslash\mathcal{A})\right)\right].$
@@
$\mathbb{E}\left[n(z)\right]=\frac{1}{4}-\frac{u^{2}(z)}{\pi^{2}}-\frac{2}{\pi^{2}}(\log(CR(z;D))-\log(CR(z;\check{D}))).$

### v2

@@ 5. Proof of the main theorem: the shifted case
Loops get a weight $2$, arcs enclosing $z$ get a weight $1\pm\alpha/2$ according to their orientation, so we obtain, using the same notations as in the preceding proof,
@@
We use the same notations as in the preceding proof: $E_{\varepsilon}$ for the zipper, $\text{sgn}(E_{\varepsilon})_{w,b}$ for the direction of an edge of the zipper. For $i\in\{1,2\}$, we write the vertices of $\mathcal{G}_{\varepsilon}^{i}$ as $(x,i)$ with $x\in\mathcal{G}_{\varepsilon}$ if $i=1$, $x\in\mathcal{G}_{\varepsilon}\setminus(\mathbb{R}\times\{0\})$ if $i=2$. Contrary to the last section, the matrix $K$ is a block matrix since $\psi=\begin{pmatrix}1\\ 0\end{pmatrix}$: $K_{(w,i),(b,j)}=0$ as soon as $i\neq j$. On the $i=j=1$ block, it is the Kasteleyn matrix on the Temperleyan domain $\mathcal{G}_{\varepsilon}^{1}$: it has Dirichlet boundary conditions everywhere. We write $K_{(w,1),(b,1)}=K^{1}_{w,b}$. On the $i=j=2$ block, it is the Kasteleyn matrix on the piecewise Temperleyan domain $\mathcal{G}_{\varepsilon}^{2}$: it has Dirichlet boundary conditions everywhere, except on the horizontal axis where it has Neumann boundary conditions (see [19] for a general discussion of mixed boundary conditions for Kasteleyn matrices). We write $K_{(w,2),(b,2)}=K^{2}_{w,b}$. The asymptotic of $(K^{1})^{-1}$ and $(K^{2})^{-1}$ when $\varepsilon\to 0$ were obtained in [19] and are recalled in Lemma C.3. Note that contrary to the last section, the vertical orientation is correct for $K^{1}$ and $K^{2}$, hence we do not need any gauge change. We can write
@@
$S_{(w,i),(b,j)}=\mathds{1}\{w\sim b\}\zeta_{w,b}\text{sgn}(E_{\varepsilon})_{w,b}(-1)^{i-1}.$ (35)
@@
$\times\left((K^{j_{i+1}})^{-1}_{b_{i},w_{i+1}}-\Delta(w_{i+1},b_{i})\right)\bigg\}.$
@@
after reindexing $r_{i+1}=r_{i}$. Now similarly to the above, if for some $i=0,\dots,n-1$, we have that $1-\kappa_{i+1}+\kappa_{i}+\tfrac{1}{2}(1-\tau_{i})$ is even (contrary to the last section), then the sum over $r_{i}=\pm 1$ gives zero. Hence the only terms that remain have for all $i=0,\dots,n-1$ that $1-\kappa_{i+1}+\kappa_{i}+\tfrac{1}{2}(1-\tau_{i})$ is odd, which is the same as $\mathds{1}\{\kappa_{i}\neq\kappa_{i+1}\}\neq\mathds{1}\{\tau=1\}$, which is the same as $\tau_{i}=(-1)^{\kappa_{i}}(-1)^{\kappa_{i+1}}$. The sum is then independent of $r$, and we obtain
@@
6. The example of a rectangle: computing the number of arcs crossing in the vertical direction
In this section, we consider a specific example of a domain - a rectangle, with $\delta$ given by two opposing sides. This is strongly inspired by section 9 of [16]. This setup doesn’t fit the precise definition of our folded or shifted dimers model, but, as in [16], we can compute new interesting quantities by directly diagonalizing the Kasteleyn matrix and using Proposition 2.1.
Let $G$ be the subgraph of $\mathbb{Z}^{2}$ with vertex set $0,n]\times[1,m]$. We define its boundary to be $\partial=\{0,n\}\times[1,m]$. We define a graph $\mathcal{G}^{r}$ by taking two copies of $G$ with vertices labeled as $(x,y,1)$ and $(x,y,2)$ for $(x,y)\in G$ and by identifying $(x,y,1)=(x,y,2)$ for $(x,y)\in\partial G$. In other words, $\mathcal{G}^{r}$ is the cylinder $\mathbb{Z}/(2n\mathbb{Z})\times[1,m]$ if we identify $(x,y,1)=(x,y)$ and (x,y,2)=(x,y)(x,y,2)=(-x,y), with the slight abuse of notation that we identify $x\in\mathbb{Z}$ with its projection $\bar{x}\in\mathbb{Z}/(2n\mathbb{Z})$ so that in particular, for all $y$, (n,y,1)=(n,y)=(n,y)=(n,y,2)(n,y,1)=(n,y)=(-n,y)=(n,y,2). A uniformly random dimer configuration on $\mathcal{G}^{r}$ gives a random configuration of disjoint loops, doubled edges and arcs, obtained by superimposing the two layers. Arcs link two boundary vertices that can be on the same side or on opposite sides of the cylinder.
@@ 6. The example of a rectangle: computing the number of arcs crossing in the vertical direction
Note that this is an example where $\partial$ has two connected components. It complements the strip from Example 3.7 where $\partial$ has one connected component.
@@
We are left with computing the left-hand side, or rather its square $\frac{\det(K_{a})}{\det{K}}$. We define the matrix $\bar{K}_{a}$ by $\bar{K}_{a}(w,b)=K_{a}(w,b)$ for all $b$ black and $w$ white in $\mathcal{G}^{r}$: it is the symmetric version of $K_{a}$ (which is by definition antisymmetric). Using the block structure of $K_{a}$ and $\bar{K}_{a}$, we get $\det(\bar{K}_{a})=-\det(K_{a})$ so the ratio of determinants is left unchanged. We now observe that the matrix $\bar{K}_{a}$ that we have constructed is exactly that of Kenyon. Writing $\lambda=\rho^{2}$ (which satisfies $a^{2n}=\lambda$, $X=\rho^{2}+\frac{1}{\rho^{2}}$), and since $K$ corresponds to $a=1$ and $X=2$, if $m$ is even, the results of Section 9 of [14] imply
@@
For $\rho\in\mathbb{C}$, such that $Y=\frac{\rho+\frac{1}{\rho}}{2}>0$, $Y^{2}=\frac{X+2}{4}$ so X>2X>-2 and all (squared) factors on both sides are non-negative so we can take the square root:
@@ Appendix A Computation in the continuum (by Avelio Sepúlveda)
The objective of this section is to compute the expected values of the random variables $n(z)$ and $o(z)$ for the ALE. These random variables are clearly conformally invariant, as the ALE itself is. It therefore suffices to compute them in a specific given domain, which we do below. Recall that the ALE has the same distribution regardless of the Dirichlet/Neumann boundary conditions of the accompanying GFF - this follows from the coupling of [18]. To be precise, we plan to show the following.
@@
Let $\mathcal{A}$ be an ALE in the strip $\mathbb{R}\times[0,\pi/2]$ with arcs emanating from the lower boundary, $n(z)$ the number of arcs of $\mathcal{A}$ that separate the upper boundary from $z$ and $O(z)=(-1)^{n(z)}$. We have that
@@
Note importantly that these moments exactly coincide with Example 3.7. Note that $n(z)$ is almost surely finite as the ALE is locally finite (any compact set intersects finitely many arcs almost surely). Unfortunately while we have formulae for all the moments of $n_{\varepsilon}(z)$ and $o_{\varepsilon}(z)$ in the $\varepsilon\to 0$ limit by Theorem 3.4, higher moments in the ALE are not currently available to us.
@@
In what follows $\lambda=\sqrt{\pi/8}$. Let $D\subseteq\mathbb{H}$ be a domain with $\partial D=\partial_{1}\cup\partial_{2}$, $\partial_{1}\cap\partial_{2}=\emptyset$ and $\partial_{2}$ an interval of $\mathbb{R}$. We also define $\check{D}$ to be the domain obtained by reflecting $D$ on $\mathbb{H}$. We take $\Phi$ to be a GFF in a domain $D$ with boundary condition $\lambda$ on $\partial_{1}$ and 0 on $\partial_{2}$, and $\check{\Phi}$ be a GFF in $D$ with boundary condition 0 on $\partial_{1}$ and free on $\partial_{2}$. We couple $\Phi$ and $\check{\Phi}$ such that $\mathbb{A}_{-\lambda,\lambda}$ of $\Phi$ coincides with $\mathcal{A}$ – the ALE of $\check{\Phi}$ as in [18] (See Figure 15). Furthermore, we denote by $u(z)$ the harmonic function in $D$ with boundary values $\lambda\mathbf{1}_{\partial_{1}}$ so that $\Phi-u(z)$ is a GFF with 0 boundary conditions in $D$.
Figure 15. Representation of $\mathbb{A}_{-\lambda,\lambda}$ and $A$ in $S^{+}$. The harmonic function associated to $\Phi_{\mathbb{A}_{-\lambda,\lambda}}$ takes alternating values in $\pm 2\lambda$, on the other hand $\check{\Phi}_{A}$ starts with value 0 on top and each time it crosses a curve changes values of $\pm 2\lambda$ independently in each connected component.
@@
Let $\Phi_{\varepsilon}$ and $\check{\Phi}_{\varepsilon}$ denote the $\varepsilon$-circle average of $\Phi$ and $\check{\Phi}$ respectively. A straightforward computation yields that (see for example Theorem 1.23 of [9])
$\mathbb{E}\left[(\Phi_{\varepsilon}(z)-u(z))^{2}\right]-\frac{1}{2\pi}\log(1/\varepsilon)\to\frac{1}{2\pi}\log(CR(z;D))\ \ \ \text{ as }\varepsilon\to 0.$ (42)
Furthermore, as $\Gamma=\frac{1}{\sqrt{2}}(\Phi-u+\check{\Phi})$ has the law of a GFF with 0 boundary conditions on $\check{D}$ restricted to $D$, we have that
$\mathbb{E}\left[(\check{\Phi}_{\varepsilon}(z))^{2}\right]-\frac{1}{2\pi}\log(1/\varepsilon)\to\frac{1}{\pi}\log(CR(z;\check{D}))-\frac{1}{2\pi}\log(CR(z;D))\ \ \ \text{ as }\varepsilon\to 0.$
@@
[…]
where $\Phi_{\mathbb{A}_{-\lambda,\lambda}}$ and $\check{\Phi}_{\mathcal{A}}$ can be represented by harmonic functions thanks to [21], and conditional on $\mathcal{A}$, $\Phi^{\mathbb{A}_{-\lambda,\lambda}}$ and $\check{\Phi}^{\mathcal{A}}$ are 0 boundary GFF on $D\backslash\mathcal{A}$ (conditionally) independent of $(\Phi_{\mathbb{A}_{-\lambda,\lambda}},\check{\Phi}_{\mathcal{A}})$. Furthermore, it is possible to check that
$\Phi_{\mathbb{A}_{-\lambda,\lambda}}(z)=\lambda O(z)\text{ and }\check{\Phi}_{\mathcal{A}}=\sum_{k=1}^{n(z)}2\lambda\xi_{k},$
@@
$u^{2}(z)=\mathbb{E}\left[(\Phi_{\varepsilon}(z))^{2}\right]=\mathbb{E}\left[((\Phi_{\mathbb{A}_{-\lambda,\lambda}})^{D}_{\varepsilon}(z))^{2}\right]+\mathbb{E}\left[((\Phi^{\mathbb{A}_{-\lambda,\lambda}})_{\varepsilon}(z))^{2}\right].$
@@
$u^{2}(z)$ $=\mathbb{E}\left[((\Phi_{\mathbb{A}_{-\lambda,\lambda}})_{\varepsilon}(z))^{2}\right]+\mathbb{E}\left[((\Phi^{\mathbb{A}_{-\lambda,\lambda}})_{\varepsilon}(z))^{2}-(\Phi_{\varepsilon}(z)-u(z))^{2}\right]$
$\to\lambda^{2}+\frac{1}{2\pi}\mathbb{E}\left[\log\left(\frac{CR(z;D\backslash\mathbb{A}_{-\lambda,\lambda})}{CR(z;D)}\right)\right]\ \ \text{ as }\varepsilon\to 0.$ (43)
@@
$\mathbb{E}[(\check{\Phi}_{\varepsilon}(z))^{2}]=\mathbb{E}\left[((\check{\Phi}_{\mathcal{A}})_{\varepsilon}(z))^{2}\right]+\mathbb{E}\left[((\check{\Phi}^{\mathcal{A}})_{\varepsilon}(z))^{2}\right].$ (44)
@@
$0$ $=\mathbb{E}\left[((\check{\Phi}_{\mathcal{A}})_{\varepsilon}(z))^{2}\right]+\mathbb{E}\left[((\check{\Phi}^{\mathcal{A}})_{\varepsilon}(z))^{2}-(\check{\Phi}_{\varepsilon}(z))^{2}\right]$
$\to\mathbb{E}\left[\left(2\lambda\sum_{k=1}^{n(z)}\xi_{k}\right)^{2}\right]+\frac{1}{2\pi}\mathbb{E}\left[-2\log(CR(z;\check{D}))+\log(CR(z;D))+\log\left(CR(z;D\backslash\mathcal{A})\right)\right]$
$=4\lambda^{2}\mathbb{E}\left[n(z)\right]+\frac{1}{2\pi}\mathbb{E}\left[-2\log(CR(z;\check{D}))+\log(CR(z;D))+\log\left(CR(z;D\backslash\mathcal{A})\right)\right].$
@@
$\mathbb{E}\left[n(z)\right]=\frac{1}{4}-\frac{2u^{2}(z)}{\pi}-\frac{2}{\pi^{2}}(\log(CR(z;D))-\log(CR(z;\check{D}))).$